OpenStudy (anonymous):

help!?!?! The volume of a cube is 36 cubic feet. What is the volume of a pyramid that fits exactly inside the cube? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (isaiah.feynman):

@luka1998 Have you seen the formulas for volume of a cube and a pyramid... to see the relation between the two shapes?

OpenStudy (anonymous):

volume of pyramid = 1/3 (B)(h) volume of cube = s^3

OpenStudy (campbell_st):

well the volume of a pyramid with the same base area as the cube will be 1/3 of the volume of the cube. so taking the known volume, 36, divide it by 3 for the volume of the pyramid

OpenStudy (campbell_st):

the pyramid will have the same base area and same height as the cube... thats why you can just divide the known volume of the cube by 3

OpenStudy (anonymous):

so the answer in word the volume of the pyramid with the same base area as the cube will be 1/3 of the volume of the cube, the volume of pyramid = the volume of the cube / 3 which is 36/3 = 12

OpenStudy (campbell_st):

correct value but remember you need units to explain the number

OpenStudy (anonymous):

what units ??

OpenStudy (isaiah.feynman):

Units of volume!

OpenStudy (campbell_st):

cubic feet...

OpenStudy (anonymous):

so the answer will be 12 cubic feet ??

OpenStudy (campbell_st):

correct.... so if you go back and re-read you'll get the explanation.

OpenStudy (anonymous):

@Isaiah.Feynman @campbell_st thx